# Aligning Translation-Specific Understanding to General Understanding in Large Language Models

### v1

@@ 5.2 Results
| COMET | QE | COMET | QE | |
| Zero-shot | 77.33 | 1.61 | 79.39 | 3.54 |
| ICL | 80.1 | 3.76 | 81.02 | 4.33 |
| xIoD-E | 81.7 | 7.00 | 81.21 | 5.75 |
The main results are illustrated in Table 1. From the results, we have drawn the following observations:
(1) xIoD achieves significant improvements.
Both xIoD-I and xIoD-E outperform the baseline ICL significantly. Specifically, xIoD-E achieves significant improvements of +1.72 and +3.31 COMET over the baseline ICL and the Zero-shot method in average, respectively. In the low-source EnIs translation, xIoD-E improves the ICL by +3.45 COMET and the Zero-shot by +5.47 COMET. These improvements demonstrate that xIoD elicits the translation ability largely through aligning the translation-specific understanding to the general one.
(2) xIoD achieves state-of-the-art performance.
xIoD achieves the highest scores in EnZh and EnEt translation in terms of COMET and achieves the highest scores across all language pairs in terms of QE. However, due to the stronger multilingual capabilities of GPT-4 over GPT-3.5, the results of xIoD based on GPT-3.5 are lower than the results of the zero-shot method based on GPT-4 in EnIs translation. To verify the effectiveness of xIoD on LLMs with stronger multilingual capabilities, we implement xIoD based on GPT-4 in EnIs translation, as shown in Table 2. The improvements achieved by xIoD suggest that LLMs with stronger multilingual capabilities also suffer from the issue of understanding misalignment.
@@
As illustrated in Table 1, CoT incurs a dramatic performance drop on the baseline ICL. To investigate this phenomenon, we have conducted case analysis and observed that the translations generated by the CoT are extremely wordy and unfluent. Our finding aligns with those reported in Peng et al. (2023). This consistency suggests that the sequentially step-by-step paradigm may not be effectively applicable to the translation task.
(4) xIoD surpasses previous translation strategies.
The results show that incorporating the analysis of keywords and topics has not achieved a consistent improvement as xIoD. We conjecture the reason is that it is the difficult-to-translate words that lead to the performance bottleneck of MT due to the long-tail distribution of knowledge Kandpal et al. (2023); Feng et al. (2023). Besides, we also follow He et al. (2023) to experiment under the rerank setting as shown in Appendix.C, which demonstrates the effectiveness of our method further.
(5) xIoD-E achieves a further improvement over xIoD-I.
Specifically, xIoD-E outperforms xIoD-I by +0.3 COMET on average. This modest improvement demonstrates that the LLM (GPT-3.5-turbo) has reached an acceptable level of detecting difficult words, and the external tool of token-level QE could enhance its capability further.
@@ 6.1 Ablation Study
xIoD introduces the processes of (1) draft translation to precisely detect the difficult words and (2) IQC to improves the correctness of interpretations. To clearly elucidate the contribution of these two components, we conduct an ablation study in Table 3. Specifically, we analyze the effect of the draft translation by asking the LLM to detect difficult words directly without the draft translation. The impact of IQC is analyzed by evaluating the performance of the generated translations guided by the original noisy interpretations (i.e., without the processing of IQC). The results show that removing either component leads to performance drops, and IQC plays a more important role in xIoD. Specifically, the improvement of xIoD is halved when ablating the IQC on the EnZh translation.
@@ 6.2 Analysis of Difficult Word Detection
To offer an in-depth insight into the process of difficult word detection, we illustrate the relation between the number of difficult words interpreted and the resulting performance by adjusting the value of the difficulty threshold (), which is shown in Fig 4. Concretely, a smaller value of allows more difficult words to be interpreted. From the results, we have the following observations:
@@
Specifically, in the EnZh translation, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to (). However, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to ().
It should be noted that interpreting more words incurs more inference costs. Therefore, a modest value of (i.e., ) is recommended to reach a compromise between efficiency and performance of xIoD.
@@ 6.3 Analysis of Interpretation Generation
Given a difficult word, xIoD generates the corresponding interpretation with the target language (i.e., cross-lingual interpretation), which implicitly comprises two stages: (1) generating the interpretation in the source language and (2) translating the interpretation into the target language. Compared with conducting these two stages explicitly, xIoD is more efficient and avoids error accumulation, which is illustrated in Fig 5. As demonstrated, interpretations in the target language (the blue bins) are more beneficial than the ones in the source language (the purple bins) owing to aligning the general understanding into the target language space, which could provide more benefits for translation. And the implicit two-stage process (the blue bins) is better than the explicit one (the green bins).
@@ 7 Related Work
With the remarkable progress of LLMs, researchers have assessed their translation abilities in various aspects. Zhang et al. (2023a); Vilar et al. (2023); Garcia et al. (2023); Bawden and Yvon (2023) have first identified best practice of LLM-based MT in terms of the prompt template and examples selection. Next, the evaluation is extended across more domains Hendy et al. (2023), more languages Zhu et al. (2023a), and document-level translation Hendy et al. (2023); Wang et al. (2023). Other lines of work have performed in-depth assessments on the important attributes beyond accuracy, like literalness Raunak et al. (2023) and culture awareness Yao et al. (2023). As existing studies have shown that LLMs have achieved promising performance especially on high-resource languages, our work turns out to benchmark them on hard instances, which are longer and more difficult to translate, towards detecting more underlying issues.
Translation-oriented fine-tuning.
To reach a new SOTA performance of MT, recent works have explored fine-tuning the LLM with parallel data. Specifically, Jiao et al. (2023a); Chen et al. (2023b) prove the effectiveness of incorporating both high-quality translations and human-annotated low-quality translations during fine-tuning. Alves et al. (2023) incorporate few-shot examples in the stage of fine-tuning to obtain both advantages of ICL and fine-tuning. Xu et al. (2023) propose the two-stage paradigm: initial fine-tuning on monolingual data followed by subsequent fine-tuning on a small set of high-quality parallel data. Zhu et al. (2023b) empower the LLM by tuning it with translation task data and cross-lingual general task data. Differently, our work aims to improve the translation performance of LLM without parameter updating, which avoids the potential risk of catastrophic forgetting on tasks beyond translation.
@@
Gao et al. (2023) introduce the domain and POS information in the inference stage. Lu et al. (2023) obtain the multilingual translations of keywords in the source sentence via the translator NLLB to augment the LLM, which improves the translation of low-resource languages while hurting the performance of high-source languages. Chen et al. (2023a) demonstrate that iterative refinement reduces translationese significantly. He et al. (2023) incorporate the knowledge of keywords, topics, and demonstrations similar to the given source sentence to enhance the translation process, respectively. Finally, a rerank-like knowledge selection is used to combine all candidate translations. However, there is no significant improvement to be observed when solely utilizing each single type of knowledge. While our xIoD could fall into the category of ”translation strategy,” we are the first to address the misalignment between the translation-specific understanding and the general understanding in LLM and achieve significant improvement in both low-resource languages and high-resource ones. In the view of methodology, different from previous works that utilize the intrinsic knowledge of LLM, xIoD focuses on dealing with the difficult-to-translate words instead of the keywords for the reason that we argue the difficult-to-translate words lead to the performance bottleneck due to the long-tail distribution of knowledge.
@@ 8 Conclusion
In this work, we take the first step in resolving the misalignment between the translation-specific understanding and the general understanding. To this effect, we propose a novel translation process xIoD, incorporating the interpretations of the difficult-to-translate words, which carry the general understanding of LLM, to guide the translation. Furthermore, we utilize the token-level QE to enhance the detection of difficult words and the sentence-level QE to remove harmful interpretations. Experimental results on the proposed Challenge-MT benchmark illustrate the effectiveness of our method.
@@ 9 Limitations
Even though xIoD elicits the translation abilities of LLMs via unleashing the general understanding (intrinsic knowledge) of LLMs, they still struggle to translate concepts that require the incorporation of extrinsic knowledge, such as the translation of neologisms. However, Our approach lays the foundation for researching when and how to incorporate external knowledge.
Appendix A Fine-grained Statistics of Challenge-MT
We compare the complete WMT test set and the Challenge-MT subset in terms of the length of source sentences, the length of target sentences, the perplexity of source sentences, average number of nouns, verbs and named entities in the source sentence. The statistics is shown in Table 4.
Appendix B Details of Experiments
@@
Appendix C Results under the Rerank setting
We follow He et al. (2023) to conduct experiment additionally under the rerank setting. For the baseline ICL, we run for 4 times with different sets of demonstrations, which are sampled randomly with seeds , and adopt QE to select the best candidate as the final translation. For MAPS, the final translation is selected from the candidates generated by the three strategies (‘+topic’, ‘+Keywords’, and ‘+SimDems’) and ICL (seed=1). For xIoD, we select the final translation from the results of xIoD and ICL (seed=1). The results are shown in Table 5.

### v2

@@ 5.2 Results on Challenge-WMT
COMET | BLEURT | COMET | BLEURT | |
Zero-shot | 77.33 | 61.15 | 79.39 | 68.40 |
ICL | 80.1 | 61.99 | 81.02 | 70.20 |
DUAT-E | 81.7 | 64.01 | 81.21 | 69.22 |
The main results are illustrated in Tab.1. From the results, we have drawn the following observations:
(1) DUAT achieves significant improvements.
On average, DUAT-E surpasses the baseline ICL by +1.82 COMET and +1.32 BLEURT, and improves Zero-shot by +3.41 COMET and +3.33 BLEURT. In the low-resource translation of EnIs, DUAT-E improves ICL by +3.85 COMET and Zero-shot by +5.87 COMET. These improvements show that, DUAT largely elicits the translation ability via aligning the translation-specific understanding to the general one.
(2) DUAT achieves state-of-the-art performance on Challenge-WMT.
DUAT achieves the highest scores in EnZh and EnEt translation in terms of COMET and BLEURT. In EnIs translation, GPT-4 achieves the best results due to its superior multilingual capabilities. To verify the effectiveness of DUAT on LLMs with stronger multilingual capabilities, we implement DUAT based on GPT-4 in EnIs translation, as shown in Table 2. The results suggest that DUAT can also benefit stronger multilingual LLMs by facilitating the understanding misalignment.
Methods | EnZh | ZhEn | ||
COMET | BLEURT | COMET | BLEURT | |
WMT22 Best | 86.80 | – | 81.00 | – |
[…]
ICL | 87.10 | 73.15 | 82.71 | 71.38 |
DUAT-E | 87.60 | 72.41 | 82.75 | 71.75 |
@@
In Table 1, CoT incurs a dramatic performance drop over the baseline ICL. Our case studies reveal that CoT produces extremely wordy translations, which is also observed by Peng et al. (2023). We conjecture that CoT makes LLMs imitate junior translators rather than advanced translators.
(4) Difficult words are the bottleneck in translating complex sentences.
The results show that incorporating the analysis of keywords and topics He et al. (2024) has yet to gain significant improvements as DUAT. It suggests that it is the difficult words that lead to the performance bottleneck in translating intricate sentences. We also follow He et al. (2024) to experiment under the rerank setting as shown in Appendix.E, which shows the effectiveness of our method further.
(5) The external tool helps LLMs better identify difficult words.
DUAT-I achieves an average improvement of +1.52 COMET, demonstrating the effective performance of LLMs in recognizing difficult words. And DUAT-E gains a further improvement of +0.3 COMET, showing the effectiveness of the external token-level QE tool in this task.
5.3 Results on the complete WMT
Experimental results on Challenge-WMT show the effectiveness of DUAT in translating complex sentences, raising the question of its impact on translating simple sentences. Therefore, we conduct additional experiments on the complete WMT2022 testset of ZhEn translation. As the results shown in Tab.5.2, our method achieves comparable results to the baseline ICL in both translation directions. These results show that DUAT has no negative impact on translating simple sentences.
EnZh | ZhEn | EnEt | EtEn | EnIs | IsEn | |
#Mistranslation | 19 | 25 | 26 | 53 | 22 | 34 |
#Generalization Failure of Baseline | 5 (26%) | 4 (16%) | 7 (27%) | 17 (32%) | 5 (23%) | 10 (29%) |
[…]
@@ 6 Analysis
6.1 Human Evaluation
To quantitatively analyze the understanding misalignment problem, we employ one senior human translator for each language direction to assess generalization failures and translation literalness. Specifically, in each direction, we randomly sample 100 sentences444It takes 1.4 dollars for annotating one sentence. from Challenge-WMT and ask the senior translator to annotate the mistranslated words and phrases (i.e., mistranslation) and score the literalness (1 to 5 score) of the translations generated by the strong baseline (GPT-3.5-turbo with ICL). Next, we ask the LLM to generate interpretations of the mistranslated content. These interpretations are presented to the translator, who judges whether they contain the correct understanding of the content. If the interpretation is accurate, the content is annotated as a generalization failure. Finally, the translator is asked to annotate the generalization failures of DUATand score the literalness of the translations produced by DUAT.
Analysis of generalization failures.
As the results shown in Tab.5.3, generalization failures account for a considerable proportion of all mistranslations (16%32%). Our method (DUAT-E) largely resolves these failures by 71%88%. We further study the unresolved failures and find that most of these unresolved failures are because they are not identified as difficult words by the LLMs in the stage of difficult word detection (Eq.3) despite actually hard to translate. It indicates that the difficult word detection remains an open question.
Analysis of translation literalness.
As the results shown in Tab.5.3, the baseline translations are highly biased towards literal translation. DUATsignificantly reduces the bias towards literal translation, indicating that the process of interpreting the difficult words first and then translating aligns better with sense-for-sense translation.
@@ 6.2 Ablation Study
DUAT introduces the processes of (1) draft translation to precisely detect the difficult words and (2) IQC to improves the helpfulness of interpretations. To clearly elucidate the contribution of these two components, we conduct an ablation study in Table 5. Specifically, we analyze the effect of the draft translation by asking the LLM to detect difficult words directly without the draft translation. The impact of IQC is analyzed by evaluating the performance of the generated translations guided by the original noisy interpretations (i.e., without the processing of IQC). The results show that removing either component leads to performance drops, and IQC plays a more important role in DUAT. Specifically, the improvement of DUAT is halved when ablating the IQC on the EnZh translation.
@@ 6.3 Analysis of Difficult Word Detection
To offer an in-depth insight into the process of difficult word detection, we illustrate the relation between the number of difficult words interpreted and the resulting performance by adjusting the value of the difficulty threshold (), which is shown in Fig.3. Concretely, a smaller value of allows more difficult words to be interpreted. From the results, we have the following observations:
@@
Specifically, in the EnZh translation, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to (). However, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to (). It should be noted that interpreting more words incurs more inference costs. Therefore, a modest value of (i.e., ) is recommended to reach a compromise between efficiency and performance of DUAT.
@@ 6.4 Analysis of Interpretation Generation
Given a difficult word, DUAT generates the corresponding interpretation with the target language (i.e., cross-lingual interpretation), which implicitly comprises two stages: (1) generating the interpretation in the source language and (2) translating the interpretation into the target language. Compared with conducting these two stages explicitly, DUAT is more efficient and avoids error accumulation, which is illustrated in Fig.4. As demonstrated, interpretations in the target language (the blue bins) are more beneficial than the ones in the source language (the purple bins) owing to aligning the general understanding into the target language space, which could provide more benefits for translation. And the implicit two-stage process (the blue bins) is better than the explicit one (the green bins).
@@ 7 Related Work
With the remarkable progress of LLMs, researchers have assessed their translation abilities in various aspects. Zhang et al. (2023a); Vilar et al. (2023); Garcia et al. (2023); Bawden and Yvon (2023) first investigate LLM-based MT in terms of the prompt template and examples selection. Next, the evaluation is extended across more domains Hendy et al. (2023), more languages Zhu et al. (2023a), and document-level translation Hendy et al. (2023); Wang et al. (2023). Other lines of work have performed in-depth assessments on the important attributes beyond accuracy, like literalness Raunak et al. (2023) and culture awareness Yao et al. (2023). As existing studies have shown that LLMs have achieved promising performance, our work turns out to benchmark them on hard instances towards detecting more underlying issues.
@@
Lu et al. (2023) obtain the multilingual translations of keywords in the source sentence via the translator NLLB to augment the LLM, which improves the translation of low-resource languages while hurting the performance of high-source languages. Chen et al. (2023) demonstrate that iterative refinement reduces translationese significantly. He et al. (2024) incorporate the knowledge of keywords, topics, and reference demonstrations to enhance the translation process, and use a rerank strategy to combine all candidate translations. However, there is no significant improvement to be observed when solely utilizing each single type of knowledge. Different from previous works that utilize the intrinsic knowledge of LLM, DUAT focuses on dealing with the difficult-to-translate words instead of the keywords for the reason that we argue the difficult-to-translate words lead to the performance bottleneck due to the long-tail distribution of knowledge.
LLM-based Automatic Post-Editing (APE).
APE corrects the errors in the generated translation, aiming to bias the translation towards the distribution of the target language (Chen et al., 2023; Koneru et al., 2024). Differently, our work aims to leverage the powerful understanding abilities of LLMs to correct the misunderstanding of complicated concepts in the source sentence. Our work is in parallel with APE.
@@ 8 Conclusion
In this work, we propose a novel translation process, DUAT, to take the first step in resolving the misalignment between the translation-specific understanding and the general understanding. Furthermore, we utilize the token-level QE tool to enhance the detection of difficult words and the sentence-level QE tool to remove harmful interpretations. Human evaluation results on high-resource and low-resource language pairs indicate that DUAT significantly facilitates the understanding alignment, which improves the translation quality and alleviates translation literalness.
@@ 9 Limitations
Even though DUAT elicits the translation abilities of LLMs via unleashing the general understanding (intrinsic knowledge) of LLMs, they still struggle to translate concepts that require the incorporation of extrinsic knowledge, such as the translation of neologisms. However, Our approach lays the foundation for researching when and how to incorporate external knowledge. Besides, DUAT requires to prompt the LLM for several times, leading to an increase in latency. This latency is mainly caused by our interpretation quality control (IQC) strategy, which sequentially ablates each generated interpretation. Concretely, if difficult words are identified, IQC needs to prompt the LLM for times.
Acknowledgements
Bing Qin is the corresponding author of this work. We thank the anonymous reviewers for their insightful comments. This work was supported by the National Natural Science Foundation of China (NSFC) (U22B2059, grant62276078), the Key R&D Program of Heilongjiang via grant 2022ZX01A32, the International Cooperation Project of PCL, PCL2022D01, the Nature Scientific Foundation of Heilongjiang Province(YQ2021F006), and the Fundamental Research Funds for the Central Universities (Grant No.HIT.OCEF.2023018).
Appendix A Generalization Failures on Translation
In this section, we first provide the illustration of generalization failures on more LLMs, as shown in Fig.5. As we can see, all of four LLMs accurately comprehend the complex concept, which three out of them mistranslate this concept. Then, we also provide more examples of generalization failures, as shown in Fig.7 and Fig.7.
Appendix B More details of DUAT
B.1 Details of IQC
We gives a formal description of our interpretation quality control in Alg.1.
B.2 Details of Demonstration Synthesis
Then, the response is parsed via regular expression to extract the difficult words and interpretations . Next, we remove the noisy interpretations through a process similar to IQC (Alg. 1). The only difference is that the QE metric is replaced with the reference-based COMET Rei et al. (2020) due to the available access to the reference translation. Finally, the generated difficult words and interpretations can be assembled with the source and target sentence as demonstrations for each step of DUAT.
Appendix C Statistics of Challenge-WMT
We compare the complete WMT test set and the Challenge-WMT subset in terms of the length of source sentences, the length of target sentences, the perplexity of source sentences, average number of nouns, verbs and named entities in the source sentence. The statistics is shown in Table 6.
Appendix D Details of Experiments
@@
Appendix E Results under the Rerank setting
We follow He et al. (2024) to conduct experiments additionally under the rerank setting. For the baseline ICL, we run for 4 times with different sets of demonstrations, which are sampled randomly with seeds , and adopt QE to select the best candidate as the final translation. For MAPS, the final translation is selected from the candidates generated by the three strategies (‘+topic’, ‘+Keywords’, and ‘+SimDems’) and ICL (seed=1). For DUAT, we select the final translation from the results of DUAT and ICL (seed=1). The results are shown in Table 7.